OpenStudy (anonymous):

Write V in the form (|v|cosΘ,|v|sinΘ), with considering the vector v=(5,8)

OpenStudy (john_es):

First, find the angle, \[\theta=\arctan(8/5)=32^o\] Then, the modulus, \[|v|=\sqrt{5^2+8^2}=\sqrt{79}\] Multiply, \[\vec{v}=(\sqrt{79}\cos\theta),\sqrt{79}\sin\theta)\]

OpenStudy (anonymous):

Wouldn't 5^2+8^2=89

OpenStudy (anonymous):

thank yo! I got it now

OpenStudy (john_es):

Yes, you are right.